### v1

# A note on restricted partition functions of Pushpa and Vasuki

Abstract.
We establish infinite families of congruences modulo arbitrary powers of 22 for three restricted partition functions $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ recently introduced by Pushpa and Vasuki by employing elementary $q$-series techniques.
Key words and phrases:
Congruences, restricted partitions, $q$-series, dissection formulas
2020 Mathematics Subject Classification:
11P83, 05A17
1. Introduction
We let $f_{m}:=\prod_{n=1}^{\infty}(1-q^{mn})$ for a positive integer $m$ and $q\in\mathbb{C}$ with $|q|<1$ throughout this paper. A partition $\lambda$ of a positive integer $n$ is a nonincreasing sequence of positive integers, known as the parts of $\lambda$, whose sum is $n$. Euler [5] showed that the generating function for the number $p(n)$ of partitions of $n$ with $p(0):=1$ is
$\sum_{n=0}^{\infty}p(n)q^{n}=\dfrac{1}{f_{1}}.$
Ramanujan [11] discovered the following remarkable congruences
$p(5n+4)$ $\equiv 0\pmod{5}$,
$p(7n+5)$ $\equiv 0\pmod{7}$,
$p(11n+6)$ $\equiv 0\pmod{11}$
for all $n\geq 0$ using classical $q$-series manipulations. Subsequently, Ramanujan [2], Watson [13], and Atkin [1] established congruences modulo powers of $5,7$, and 1111 for $p(n)$.
Pushpa and Vasuki [9] introduced in 2022 seven restricted partition functions, three of which have the following generating functions defined by
$\sum_{n=0}^{\infty}M(n)q^{n}$ $:=\dfrac{f_{2}^{5}f_{5}^{5}}{f_{1}f_{10}},$
$\sum_{n=0}^{\infty}T^{\ast}(n)q^{n}$ $:=\dfrac{f_{1}^{5}f_{10}^{5}}{f_{2}f_{5}},$
$\sum_{n=0}^{\infty}P^{\ast}(n)q^{n}$ $:=f_{1}^{4}f_{5}^{4},$
and proved some congruences modulo 22 and 55 for $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$. Consequently, Dasappa et. al. [4] obtained more congruences and linear identities for these restricted partition functions. Very recently, Nath and Saikia [8] established infinite families of congruences modulo powers of 55 for $M(n)$ and $T^{\ast}(n)$ and particular congruences for $P^{\ast}(n)$, and settled the conjecture of Dasappa et. al. [4, Conjecture 1] by relying on the Mathematica package RaduRK created by Smoot [12], which is based on Radu’s Ramanujan–Kolberg algorithm [10].
In this short note, we explore more arithmetic properties of $M(n),T^{\ast}(n)$, and $P^{\ast}(n)$ using elementary $q$-series techniques. Our main results reveal infinite families of congruences modulo powers of 22 for these restricted partition functions, which generalize some results of Nath and Saikia [8].
Theorem $1.$
For all integers $k\geq 1$ and $n\geq-1$, we have
$M(2^{k}n+2^{k+1}-1)$ $\equiv 0\pmod{2^{k-1}}$, (1.1)
$T^{\ast}(2^{k}n+2^{k+1}-2)$ $\equiv 0\pmod{2^{k-1}}$. (1.2)
Theorem 1.2.
For all integers $k\geq 0$ and $n\geq-1$, we have
$P^{\ast}(2^{4k}n+2^{4k+1}-1)$ $\equiv 0\pmod{2^{6k}}$, (1.3)
$P^{\ast}(2^{4k+1}n+2^{4k+2}-1)$ $\equiv 0\pmod{2^{6k+2}}$, (1.4)
$P^{\ast}(2^{4k+2}n+2^{4k+3}-1)$ $\equiv 0\pmod{2^{6k+3}}$, (1.5)
$P^{\ast}(2^{4k+3}n+2^{4k+4}-1)$ $\equiv 0\pmod{2^{6k+6}}$, (1.6)
$P^{\ast}(2^{4k+4}n+3\cdot 2^{4k+3}-1)$ $=0.$ (1.7)
We organize the rest of the paper as follows. In Section 2, we present some theta function identities needed to prove Theorems 1.1 and 1.2, including the generating functions for $M(2n+3),T^{\ast}(2n+2)$, and $P^{\ast}(2n+3)$. In Section 3, we demonstrate Theorems 1.1 and 1.2 by finding the generating functions for $M(2^{k}n+2^{k+1}-1),T^{\ast}(2^{k}n+2^{k+1}-2)$, and $P^{\ast}(2^{k}n+2^{k+1}-1)$ for any integer $k\geq 1$ via induction.
2. Preliminaries
We provide in this section necessary theta function identities to prove Theorems 1.1 and 1.2.
Lemma 2.1.
We have the identities
$\dfrac{f_{2}^{5}}{f_{10}}$ $=\dfrac{f_{1}^{5}}{f_{5}}+5q\dfrac{f_{1}^{2}f_{2}f_{10}^{3}}{f_{5}^{2}},$ (2.1)
$\dfrac{f_{1}}{f_{5}}$ $=\dfrac{f_{2}f_{8}f_{20}^{3}}{f_{4}f_{10}^{3}f_{40}}-q\dfrac{f_{4}^{2}f_{40}}{f_{8}f_{10}^{2}},$ (2.2)
$f_{1}f_{5}^{3}$ $=f_{2}^{3}f_{10}-q\dfrac{f_{2}f_{8}^{2}f_{20}^{6}}{f_{4}^{2}f_{10}f_{40}^{2}}+2q^{2}f_{4}f_{20}^{3}-q^{3}\dfrac{f_{4}^{4}f_{10}f_{40}^{2}}{f_{2}f_{8}^{2}}.$ (2.3)
Proof.
From [3, Theorem 10.4], we have
$\dfrac{f_{2}f_{5}^{5}}{qf_{1}f_{10}^{5}}$ $=\dfrac{1}{k(q)}-k(q),$ (2.4)
$\dfrac{f_{2}^{4}f_{5}^{2}}{qf_{1}^{2}f_{10}^{4}}$ $=\dfrac{1}{k(q)}+1-k(q),$ (2.5)
$\dfrac{f_{1}^{3}f_{5}}{qf_{2}f_{10}^{3}}$ $=\dfrac{1}{k(q)}-4-k(q).$ (2.6)
where
$k(q):=q\prod_{n=1}^{\infty}\dfrac{(1-q^{10n-9})(1-q^{10n-8})(1-q^{10n-2})(1-q^{10n-1})}{(1-q^{10n-7})(1-q^{10n-6})(1-q^{10n-4})(1-q^{10n-3})}$
is the Ramanujan’s function. Subtracting (2.6) from (2.5) leads to
$\dfrac{f_{2}^{4}f_{5}^{2}}{qf_{1}^{2}f_{10}^{4}}-\dfrac{f_{1}^{3}f_{5}}{qf_{2}f_{10}^{3}}=5.$ (2.7)
Multiplying both sides of (2.7) by $qf_{1}^{2}f_{2}f_{10}^{3}/f_{5}^{2}$, we obtain (2.1). Replacing $q$ with q-q in [7, Theorem 2.1] and using the fact that
$\prod_{n=1}^{\infty}(1-(-q)^{n})=\dfrac{f_{2}^{3}}{f_{1}f_{4}},$
we obtain (2.2). We next infer from (2.4) and (2.5) that
$\dfrac{f_{2}f_{5}^{5}}{qf_{1}f_{10}^{5}}=\dfrac{f_{2}^{4}f_{5}^{2}}{qf_{1}^{2}f_{10}^{4}}-1.$ (2.8)
Multiplying both sides of (2.8) by $qf_{1}^{2}f_{10}^{5}/(f_{2}f_{5}^{2})$ and using (2.2), we see that
$f_{1}f_{5}^{3}$ $=f_{2}^{3}f_{10}-q\dfrac{f_{1}^{2}f_{10}^{5}}{f_{2}f_{5}^{2}}=f_{2}^{3}f_{10}-q\dfrac{f_{10}^{5}}{f_{2}}\left(\dfrac{f_{2}f_{8}f_{20}^{3}}{f_{4}f_{10}^{3}f_{40}}-q\dfrac{f_{4}^{2}f_{40}}{f_{8}f_{10}^{2}}\right)^{2}.$ (2.9)
Expanding (2.9) yields (2.3). ∎
Lemma $2.$
Let $q^{-3}f_{1}^{4}f_{5}^{4}=\sum_{n=-3}^{\infty}P^{\ast}(n+3)q^{n}$. Then
$\sum_{n=-1}^{\infty}P^{\ast}(2n+3)q^{n}=-4\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8f_{2}^{4}f_{10}^{4}.$
Proof.
See [6, Proposition 3.2]. ∎
We now apply Lemmas 2.1 and 2.2 to derive the generating functions for $M(2n+3)$ and $T^{\ast}(2n+2)$.
Lemma 2.3.
We have that
$\sum_{n=-1}^{\infty}M(2n+3)q^{n}$ $=\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8f_{2}^{4}f_{10}^{4}+10f_{1}f_{2}f_{5}^{3}f_{10}^{3},$ (2.10)
$\sum_{n=-1}^{\infty}T^{\ast}(2n+2)q^{n}$ $=\dfrac{f_{1}^{4}f_{5}^{4}}{q}+10f_{1}f_{2}f_{5}^{3}f_{10}^{3}.$ (2.11)
Proof.
Using (2.1) and (2.3), we have
$\sum_{n=-3}^{\infty}M(n+3)q^{n}$ $=\dfrac{f_{2}^{5}f_{5}^{5}}{q^{3}f_{1}f_{10}}=\dfrac{f_{5}^{5}}{q^{3}f_{1}}\left(\dfrac{f_{1}^{5}}{f_{5}}+5q\dfrac{f_{1}^{2}f_{2}f_{10}^{3}}{f_{5}^{2}}\right)$
$=\dfrac{f_{1}^{4}f_{5}^{4}}{q^{3}}+5\dfrac{f_{1}f_{2}f_{5}^{3}f_{10}^{3}}{q^{2}}$
$=\dfrac{f_{1}^{4}f_{5}^{4}}{q^{3}}+5\dfrac{f_{2}f_{10}^{3}}{q^{2}}\left(f_{2}^{3}f_{10}-q\dfrac{f_{2}f_{8}^{2}f_{20}^{6}}{f_{4}^{2}f_{10}f_{40}^{2}}+2q^{2}f_{4}f_{20}^{3}-q^{3}\dfrac{f_{4}^{4}f_{10}f_{40}^{2}}{f_{2}f_{8}^{2}}\right).$ (2.12)
We consider the terms involving $q^{2n}$ in the expansion of (2). In view of Lemma 2.2, we have
$\sum_{n=-1}^{\infty}M(2n+3)q^{n}$ $=-4\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8f_{2}^{4}f_{10}^{4}+5\dfrac{f_{1}^{4}f_{5}^{4}}{q}+10f_{1}f_{2}f_{5}^{3}f_{10}^{3},$
which is exactly (2.10).
On the other hand, we apply (2.1) and (2.2) so that
$\sum_{n=-2}^{\infty}T^{\ast}(n+2)q^{n}$ $=\dfrac{f_{1}^{5}f_{10}^{5}}{q^{2}f_{2}f_{5}}=\dfrac{f_{10}^{5}}{q^{2}f_{2}}\left(\dfrac{f_{2}^{5}}{f_{10}}-5q\dfrac{f_{1}^{2}f_{2}f_{10}^{3}}{f_{5}^{2}}\right)$
$=\dfrac{f_{2}^{4}f_{10}^{4}}{q^{2}}-5\dfrac{f_{1}^{2}f_{10}^{8}}{qf_{5}^{2}}=\dfrac{f_{2}^{4}f_{10}^{4}}{q^{2}}-5\dfrac{f_{10}^{8}}{q}\left(\dfrac{f_{2}f_{8}f_{20}^{3}}{f_{4}f_{10}^{3}f_{40}}-q\dfrac{f_{4}^{2}f_{40}}{f_{8}f_{10}^{2}}\right)^{2}.$ (2.13)
Extracting the terms involving $q^{2n}$ in the expansion of (2) yields (2.11). ∎
3. Proofs of Theorems 1.1 and 1.2
We establish in this section Theorems 1.1 and 1.2 by using theta function identities and certain generating functions derived from Section 2. As an application of these identities, we derive the following generating functions which are instrumental in the proofs of the aforementioned theorems.
Theorem 3.1.
For all $k\geq 1$, we have that
$\sum_{n=-1}^{\infty}M(2^{k}n+2^{k+1}-1)q^{n}$ $=A_{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8A_{k-1}f_{2}^{4}f_{10}^{4}+5\cdot 2^{k}f_{1}f_{2}f_{5}^{3}f_{10}^{3},$ (3.1)
$\sum_{n=-1}^{\infty}T^{\ast}(2^{k}n+2^{k+1}-2)q^{n}$ $=B_{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8B_{k-1}f_{2}^{4}f_{10}^{4}+5\cdot 2^{k}f_{1}f_{2}f_{5}^{3}f_{10}^{3},$ (3.2)
$\sum_{n=-1}^{\infty}P^{\ast}(2^{k}n+2^{k+1}-1)q^{n}$ $=C_{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8C_{k-1}f_{2}^{4}f_{10}^{4},$ (3.3)
where for $k\geq 2$, $P_{k}\in\{A_{k},B_{k}\}$ and $C_{k}$ satisfy the recurrence
$P_{k}$ $=-4P_{k-1}-8P_{k-2}+5\cdot 2^{k-1},$
$C_{k}$ $=-4C_{k-1}-8C_{k-2},$
with initial values $B_{0}=0,A_{0}=A_{1}=B_{1}=C_{0}=1$, and $C_{1}=-4$.
Proof.
We only prove (3.1), as the proofs of (3.2) and (3.3) are similar to that of (3.1). We proceed by induction on $k$. In view of Lemmas 2.2 and 2.3, we note that (3.1)–(3.3) hold for $k=1$. Suppose (3.1) holds for some $k\geq 2$. Dividing both sides of (3.1) by $q^{2}$ and employing (2.3), we find that
$\sum_{n=-3}^{\infty}$ $M(2^{k}n+2^{k+2}-1)q^{n}$
$=A_{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q^{3}}-8A_{k-1}\dfrac{f_{2}^{4}f_{10}^{4}}{q^{2}}+5\cdot 2^{k}\dfrac{f_{1}f_{2}f_{5}^{3}f_{10}^{3}}{q^{2}}$
$=A_{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q^{3}}-8A_{k-1}\dfrac{f_{2}^{4}f_{10}^{4}}{q^{2}}+5\cdot 2^{k}\dfrac{f_{2}f_{10}^{3}}{q^{2}}\left(f_{2}^{3}f_{10}-q\dfrac{f_{2}f_{8}^{2}f_{20}^{6}}{f_{4}^{2}f_{10}f_{40}^{2}}+2q^{2}f_{4}f_{20}^{3}-q^{3}\dfrac{f_{4}^{4}f_{10}f_{40}^{2}}{f_{2}f_{8}^{2}}\right).$ (3.4)
We read the terms involving $q^{2n}$ in the expansion of (3). We see from Lemma 2.2 that
$\sum_{n=-1}^{\infty}$ $M(2^{k+1}n+2^{k+2}-1)q^{n}$
$=A_{k}\left(-4\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8f_{2}^{4}f_{10}^{4}\right)-8A_{k-1}\dfrac{f_{1}^{4}f_{5}^{4}}{q}+5\cdot 2^{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}+5\cdot 2^{k+1}f_{1}f_{2}f_{5}^{3}f_{10}^{3}$
$=(-4A_{k}-8A_{k-1}+5\cdot 2^{k})\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8A_{k}f_{2}^{4}f_{10}^{4}+5\cdot 2^{k+1}f_{1}f_{2}f_{5}^{3}f_{10}^{3}$
$=A_{k+1}\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8A_{k}f_{2}^{4}f_{10}^{4}+5\cdot 2^{k+1}f_{1}f_{2}f_{5}^{3}f_{10}^{3},$
where the last equality follows from the definition of $A_{k}$. Thus, (3.1) holds for $k+1$, so (3.1) holds for all $k\geq 1$ by induction. ∎
Proof of Theorem $1.$
We claim that for all $k\geq 1$ and $P_{k}\in\{A_{k},B_{k}\}$, $P_{k}$ is divisible by $2^{k-1}$ but not by $2^{k}$. We prove this via strong induction on $k$. We compute $A_{2}=6$ and $B_{2}=-2$, so the claim holds for $k\in\{1,2\}$. Suppose now that the claim holds for all positive integers at most some $k\geq 3$. Write $P_{k}=2^{k-1}a_{k}$ and $P_{k-1}=2^{k-2}b_{k}$ for some odd integers $a_{k}$ and $b_{k}$. We infer from the definition of $P_{k}$ that
$P_{k+1}=-4P_{k}-8P_{k-1}+5\cdot 2^{k}=2^{k}(-2a_{k}-2b_{k}+5).$
Since $5-2a_{k}-2b_{k}$ is odd, we obtain $2^{k}\mid P_{k}$ but $2^{k+1}\nmid P_{k}$. Thus, the claim holds for $k+1$, and hence for $k\geq 1$ by strong induction. Combining the above claim and the generating functions (3.1) and (3.2), we arrive at (1.1) and (1.2), respectively. ∎
To prove Theorem 1.2, we require the following result on the exact values of $C_{k}$.
Lemma 3.2.
For all integers $k\geq 0$, we have
$C_{4k}$ $=(-64)^{k},$
$C_{4k+1}$ $=-4(-64)^{k},$
$C_{4k+2}$ $=8(-64)^{k},$
$C_{4k+3}$ $=0$
Proof.
Using the definition of $C_{k}$, we observe that for all $k\geq 0$ and $r\in\{0,1,2,3\}$,
$C_{4k+r+4}$ $+64C_{4k+r}$
$=C_{4k+r+4}+4C_{4k+r+3}+8C_{4k+r+2}-4(C_{4k+r+3}+4C_{4k+r+2}+8C_{4k+r+1})$
$+8(C_{4k+r+2}+4C_{4k+r+1}+8C_{4k+r})=0.$ (3.5)
We see from (3.5) that $C_{4k+r}=(-64)^{k}C_{r}$ for all $k\geq 0$ and $r\in\{0,1,2,3\}$. In view of the computed values $C_{2}=8$ and $C_{3}=0$, we arrive at the desired values of $C_{k}$. ∎
Proof of Theorem 1.2.
Appealing to Lemma 3.2, we first replace $k$ with $4k$ in (3.3) so that
$\sum_{n=-1}^{\infty}P^{\ast}(2^{4k}n+2^{4k+1}-1)q^{n}$ $=(-64)^{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q},$
yielding (1.3). We next replace $k$ with $4k+1$ in (3.3) so that
$\sum_{n=-1}^{\infty}P^{\ast}(2^{4k+1}n+2^{4k+2}-1)q^{n}$ $=4(-64)^{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}-8(-64)^{k}f_{2}^{4}f_{10}^{4},$
which gives (1.4). We next replace $k$ with $4k+2$ in (3.3) so that
$\sum_{n=-1}^{\infty}P^{\ast}(2^{4k+2}n+2^{4k+3}-1)q^{n}$ $=8(-64)^{k}\dfrac{f_{1}^{4}f_{5}^{4}}{q}+32(-64)^{k}f_{2}^{4}f_{10}^{4},$
which gives (1.5). We now replace $k$ with $4k+3$ in (3.3) so that
$\sum_{n=-1}^{\infty}P^{\ast}(2^{4k+3}n+2^{4k+4}-1)q^{n}$ $=(-64)^{k+1}f_{2}^{4}f_{10}^{4},$ (3.6)
arriving at (1.6). Dividing both sides of (3.6) by $q$, we have
$\sum_{n=-2}^{\infty}P^{\ast}(2^{4k+3}n+3\cdot 2^{4k+3}-1)q^{n}$ $=(-64)^{k+1}\dfrac{f_{2}^{4}f_{10}^{4}}{q}.$ (3.7)
Examining the terms involving $q^{2n}$ in the expansion of (3.7), we finally obtain (1.7). ∎